OpenStudy (anonymous):

HELP PLEASE!!! Evaluate the following expression using the values given: Find -a^2 - 3b^3 + c^2 + 2b^3 - c^2 if a = 3, b = 2, and c = -3. -67 -17 -31 1

OpenStudy (anonymous):

just plug in a b and c into the equation and u should find your answer

OpenStudy (anonymous):

yes i know that i did that and i keep coming up with the same answer and it was wrong

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

after you make the proper substitutions, you should have this -a^2 - 3b^3 + c^2 + 2b^3 - c^2 -(3)^2 - 3(2)^3 + (-3)^2 + 2(2)^3 - (-3)^2

jimthompson5910 (jim_thompson5910):

You would then evaluate -(3)^2 - 3(2)^3 + (-3)^2 + 2(2)^3 - (-3)^2 using PEMDAS

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

9 - 24 + 9 + 16 - 9 = 1 which is the wrong answer

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

you made an error on the first term -(3)^2 turns into -9 and not +9

jimthompson5910 (jim_thompson5910):

you should have this -9 - 24 + 9 + 16 - 9

OpenStudy (anonymous):

how did u get -9

jimthompson5910 (jim_thompson5910):

at the very left is -a^2 plug in a = 3 -a^2 = -(3)^2 = -9

OpenStudy (anonymous):

OOOOOOOOOOOOOOOOOOhhhhhhhhh now i see my mistake

OpenStudy (anonymous):

THANK YOU YOU THE BEST!!! :D

jimthompson5910 (jim_thompson5910):

you're welcome